OpenStudy (moazzam07):

The expression 5x − 3(2x − 4) is equivalent to the expression 12 − x.

OpenStudy (moazzam07):

true or False

OpenStudy (horsegirl27):

wrong section:P

OpenStudy (moazzam07):

what?

OpenStudy (horsegirl27):

You're in OpenStudy Feedback. Please post in the mathematics section.

OpenStudy (anonymous):

= 5x − 3(2x − 4) = 5x - 6x + 12 = 12 - x This is true dear Best of luck